LHC Machine Outreach

Blog 2008

Photo Gallery

Friday 21st December

Last day at work before CERN shuts down for the Christmas break. The cryogenics group have got a very good handle on sector 45 and the main dipole string was taken to 8500 Amps on Wednesday (equivalent to around 5 TeV in beam energy).

The cool down of sector 56 continues and it is around 100 K today.

LHC visible in Google Earth

Friday 7th December

Nice set of LHC pictures in today's Guardian

Things have been a bit bumpy in sector 45, with the cryogenics group wrestling hard with a number of problems. Despite this magnet commissioning has started with an insertion quadrupole powered to nominal yesterday. Sector 56 has started it's descent to low temperature.

Saturday 1st December

Sector 45 is cold - the second of the eight LHC sectors to get down to around 1.9 K. Powering tests will start in the first week of December. The third sector, 56, has started the cool down process.

Friday 30th November 2007

Right in time for the festive season, ESA's XMM-Newton X-ray observatory has discovered a huge cloud of high-temperature gas resting in a spectacular nearby star-forming region, shaped somewhat like the silhouette of Santa Claus.


Saturday 27th October

String theory in 2 minutes or less!

Wednesday 24th October

Cool down of sector 45 progressing well.

Monday 8th October

Scientists Ask Congress To Fund $50 Billion Science Thing

(Not that we don't take ourselves very seriously here at CERN.)

Thursday 4th October 2007

Joining of all the magnets up is effectively completed, and attention has switched to testing and cooling everything down.

Understandably, given the complex nature of the beast there have been some inevitable problems, and key among these is the appearance of vacuum leaks during the cool down procedure. The drop from room temperature to 1.9 degrees above absolute zero causes considerable contraction and the stress induced will expose any weakness.

Lately this has caused delays in the cool down of sector 4-5 (the second of the eight sectors to cooled). Two vacuum leaks caused the process to be halted - locating the leaks is awkward and usually the sector has to be warmed up to fix the problem.

In sector 8-1, another leak forced the replacement of a quadrupole with inevitable delay in the cool down of that sector. However, the vacuum group (a high skilled and motivated team!) are working hard to resolve these issues and its hope that good progress in cooling four to five sectors will have been made by the end of the year.

Another recent problem involves the so-called plug-in modules which link the beam vacuum tubes between magnets. To allow contraction and expansion of the magnets during cool down and warm-up these elements have sliding parts which move over each other (see pictures below). Unfortunately a small number of these in sector 78 have failed to slide back during the warm-up and parts have buckled into the beam pipe - not good at all when it comes to getting the beam around. Happily the problem has been spotted and diagnostic techniques have been developed to spot the faulty components - these include x-ray and blowing a small ball with a transmitter inside around the ring.

Normal plug in module

Failed module - notice the buckled fingers projected into the beam aperture

Tuesday August 14th 2007

Latest possibility for things that might be found at the LHC: unparticles

Monday August 6th 2007

[The LHC is divided into 8 more-or-less independent sectors - each of these sectors have to be cooled and commissioned in turn before we can take beam.]

Sector 78 is being warmed up after a successful cool-down and magnet tests. Sector 45 should be in cool-down but this is stalled because of a vacuum leak. Details below. These sorts of problems are almost inevitable for a machine of this complexity.

Sector 45
After the detection and the localization of a leak in the vacuum sector closest to Point 4, the cool down of Sector 45 was stopped and the active warm up of four cells started last Wednesday July 18th: on Monday July 23rd the insulation vacuum was vented to atmospheric pressure with dry air and four interconnections were opened on Tuesday. Measurements indicated that the leak is localized on the DFBAH side. On Thursday July 26th, the front plate of the HCM of the DFBAH was opened and the measurement was confirmed. It is suspected that the leak is in the jumper connection on the HCM module of the DFBAH: the jumper connection is presently being opened. The rest of the sector is left floating in temperature.

Sector 78
The warm-up of Sector 78 was stopped on Saturday July 22nd at noon after the short circuit on the RQD circuit which had been observed in the course of the cool down reappeared. On Monday July 23rd the ElQA team confirmed the localization of the short circuit and the warm-up on restarted on Tuesday morning. The sector is expected to be warm on August 6th and ready for the connection of the inner triplet assembly left of Point 8 and consolidation work.

Sector 81
The work to localize the vacuum leak which was observed after the pressure tests in Sector 81 continues. The cool down is expected to start on week 36. In parallel with this, an additional campaign of the warm magnets which aims at increasing the noise rejection of the power converters around Point 8 will be completed this Friday July 27th. The electrical quality assurance of the closed orbit corrector circuits at warm in the arc is progressing.

Sunday July 22nd 2007

Sector 78 after an encouraging hardware commissioining is now being warmed up - need to make some repairs and connect the inner triplets before a re-cool down later this year..

Sector 45 is being cooled.

The repair of the inner triplets has been implemented on an installed assembly and looks good - all installed assemblies will now be fixed and tested in situ.

It's summer. The car parks are half empty. Summer students are partying hard and the place seems full of Americans - presumably because they don't get as much holiday as us Europeans.

Wednesday 27th June 2007

Main dipole string in sector 78 up to 2000A.

Saturday 23rd June 2007

Slippage of the schedule announced.
CERN announces new start-up schedule for world’s most powerful particle accelerator

Inevitable delays have been accumulating, not helped by the inner triplet problems.

However, the backdrop to this is the very good behaviour of the cryogenics in sector 7-8 (see hardware commissioning progress link on the right) and the ongoing commissioning of the magnets in that region. A couple of days ago the main defocusing quadrupole magnet chain was brought up to 6500A and the main dipoles are next up. In the case of the dipoles this means powering all 154 14.3 m. dipole magnets which bend the beam around the 3 km. arc. This really represents the bread and butter of LHC operations and getting these things on-line would certainly be an important physiological step. Lots to do before we get the whole ring cold and tested, but mood is optimistic.

Friday 25th May

There's been a bit of excitement in the LHC control room this week.  For the first time some of the high power circuits of sector 7-8 have been powered.   For  the moment it is the insertion quadrupoles, Q4 and Q5, close to point 8,  that have been worked on. There have been several quenches (both provoked and training) and lots of data generated for analysis.  All these elements are operated at 4.5K.   In addition powering of the low current correctors in the arc and in the insertions is continuing.

Next week, there is now a good hope that powering of the main circuits in arc 7-8 will finally begin.   Many problems in the cryogenic system have been identified and most corrected.    The cryogenic system now seems much more stable and cool-down of the arc to 1.9K is progressing well.  The average temperature in the arc is now around 2.3K.

Friday 27th April 2007

Last dipole down      The first dipole went down on 7th March 2005.

Friday 13th April 2007

Sector 78 has reached 1.9 K   Forget the inner triplet problems - this is a major achievement.

Wednesday 4th April 2007

The cryogenic team are getting very close - some magnets in 7-8 are already at superfluid helium temperatures.

magnet temperatures along arc

Friday 30th March 2007

The big news this week is the failure of the inner triplet assembly left of point 5 during a pressure test.
Fermilab Statement on LHC Magnet Test Failure

The inner triplets consists of 4 strong superconducting quadrupoles which are used (with other nearby quadrupoles) to focus the beam down at the interaction point inside the neighbouring experiment (in this case CMS).

Thursday 22nd March


Thursday 8th March

Day to Day Communications at CERN in 1974
(La plus ça change...)

and the magnets in 7-8 are getting colder:

Wednesday 21st February

A hell of lot going on out there! A very brief overview...

Cool-down of Sector 7-8 started mid-January
• Interconnection on-going in 5 sectors:
o Start of sector 6-7 mid-February
o Closure of sector 4-5 end of February
• 16 DFBM (70%), 8 DFBA (50%) and 1 DFBL (20%) in place (The DFBs are the electrical feedboxes which bring the current from room temperature into the cold mass.)
• (4 DFBM available; not installed - delays on DSLC or corresponding SSSS)
• Rate of warm beam-pipe installation is ramping up:
o Coordination with individual system tests is tight
• 1542 cryo-magnets in place (88%)

Cooldown of the first sector continues...


Wednesday 31st January

The status of the cooldown is shown below. An earth fault on one of the main quadrupole circuits did develop last week (electical quality assurance tests are performed periodically on the way down) but fortunately this appears have disappeared as the temperature was lower further.


Thursday 25th January

In the plot below you can see the progress of the cooldown of sector 7-8. The green line shows the magnet temperature in Kelvin (K). The aim: 1.9 K. We've got snow on the ground here in Geneva at last, and it's interesting to note that the cooldown was slowed down because it was too cold outside. The lorries carrying the liquid nitrogen couldn't get through.

Monday 15th January 2007

The cooldown of the first of the eight of the LHC sectors started today.